INAF queue observing, 2016 May 03-04

Observer: R. Carini, M. Di Criscienzo
Telescope Operator: D. Gonzalo-Huerta
Support Astronomer: O. Kuhn
Instrument: LBC, MODS

Summary

At the sunset the sky is clear. At the midnight the wind grows up At 8.15 the TO closes the dome

Night duration (12\xB0 el twilight) 20:30-4:30 ( 03:30-11.30 UT) = 8 h 00m

Observing time: * 4h 45m* (59%)

Weather time loss: * 3h 15m* (41%)

technical time loss * h* (00%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
FLAT LBC   r, R, I in the red channel    
Giallongo Virgo field 2 dual grating, r sloan, 2h 15m the ob in LBC virgo_rsloan_180s_90deg.ob is not completed
Giallongo Virgo field 1 dual grating, r sloan 1 h the blue images are not good, in the red the spectra are visible
DARK LUCI        
MODS CALIB flat and arc for standard, bias dual grating    

Details

02:00 Sky Flat for LBCred:
  • images in r slon are good at 180deg
  • images in R are good in 180deg and 1 in 0deg
  • two images in I are good at 0 deg

03:05 acquisition standard star for Giallongo's program
  • acqMODS feige67.acq mods1r.20160504.0001-0002
  • execMODS feige67_ls.obs mods1r.0003, mods1b.20160504.0001
  • seeing 1.0''

03:30 slew to Giallongo's program
  • acqMods virgo2_mos.acq mods1r.0004-0006
  • 3.45 execMODS virgo2_mos_1.obs mods1r.0007-0008, mods1b.0002-0003
  • 03:50 run virgo_rsloan_10s_90deg.ob
  • 03:58 run virgo_rsloan_10s_180deg.ob
  • 04:07 virgo_rsloan_10s_270deg.ob
  • 04:16 virgo_rsloan_60s_90deg.ob
  • seeing about 1.0''

04:25 execMODS virgo2_mos_2.obs mods1r.0009-0012, mods1b.0004-0007
  • 04:38 dofpia
  • 04.45 run virgo_rsloan_60s_180deg.ob * seeing 1.0'' * 04.55 seeing about 1.2'' * 05:03 virgo_rsloan_60s_270deg.ob * 05:06 seeing 0.8''-0.1'' * 05:21 dofpia * 05:26 virgo_rsloan_180s_0deg.ob * seeing about 0.8'''

05:34 execMODS virgo2_mos_3.obs mods1r.0013-0015, mods1b.0008-0010
  • 06:10 virgo_rsloan_180s_90deg.ob
  • 06:30 the target is out of spec (for the parallatic angle), we stop the acquisition, in LBC the images are 5/10, but the last one is not good (062521.fits) finished the script with STOP we have save also the images 0016 in red and 0011 in blue but are not good

06:30 slew to field Virgo1
  • acqMODS virgo1_mos.acq mods1r.0017-0020
  • dofpia
  • during the acquisition the wind increases suddenly, seeing increases with peaks to 1.9''
  • execMODS virgo1_mos_2.obs mods1r.0021-0024, mods1b.0012-0015 the blue images are non good, in the red the spectra are visible
  • the seeing is very variable 1.4''-1,6''
  • 07:06* seeing variable, Avg FWHM =1.5
  • 07:09 run virgo_rsloan_10s_90deg.ob
  • 07:13 seeing 1.3
  • 07:18 virgo_rsloan_10s_180deg.ob
  • 07:20 seeing about 1.5'', wind velocity 18.4 m/s
  • 07:32 seeing 1.3''-1.5'', wind velocity 17.1 m/s
  • 07:37 virgo_rsloan_10s_270deg.ob
  • 07:45 virgo_rsloan_60s_90deg.ob (3 ima)
  • 07:50 dofpia and again run virgo_rsloan_60s_90deg.ob

08:18 wind rear velocity 20.1. The TO closes the dome. we have done other 6 images with LBC

08:27 DARK LUCI*
  • dark for the target A4 in H2 luci2.20160504.0003-0012
  • dark for the target A4 in Ks luci2.20160504.0013-0022
  • dark for the standard star luci2.20160504.0023-0032

09:38 MODS CALIBRATION
  • flat with slit 0.5'' execMODS slitflat_5as.cal mods1r.0025-0027, mods1b.0016-0021
  • arc with slit 0.5''' execMODS grlamps.cal mods1r.0028-0030, mods1b.0022-0024
  • bias mods1r.0031-0035, mods1b.0025-0029

-- %USERSIG{RobertaCarini - 2016-05-04}%

Comments

 
Topic revision: r6 - 06 May 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback